OpenStudy (anonymous):

Solve the system, using substitution. y = 1 - 5x and 5x + y = 8 A. no solution B. infinitely many solutions

OpenStudy (anonymous):

@ganeshie8

jimthompson5910 (jim_thompson5910):

5x + y = 8 5x + 1-5x = 8 ... replace the 'y' with '1-5x' (since we know y = 1-5x) 0x + 1 = 8 0 + 1 = 8 1 = 8 ... this is ALWAYS false regardless of what x and y are. So there are NO solutions

jimthompson5910 (jim_thompson5910):

y = 1 - 5x and 5x + y = 8 are parallel lines that never cross. You need an intersection point to have a solution